The wage rate of labor is Rs. 6 and price of capital is Rs. 2. The marginal product of labor is 16 while marginal product of capital is 4. Can a firm be operating at equilibrium?     


1
Expert's answer
2021-05-27T10:58:08-0400

The formula for the firm's equilibrium is ;"\frac {MP_L} {w} =\frac {MP_K} {r}"


By substituting these values in the equilibrium condition ;


"\frac {16} {6} =\frac {4} {2}"


"\frac {8} {3} =\frac {2} {1}"


"2.67\gt2"


This demonstrates that the provided input combination is inefficient, as the firm may produce more output by spending an additional unit of money on labor rather than capital. As a result, the firm is not in equilibrium, and in order to operate efficiently in equilibrium, the business needs substitute labor for capital.